OpenStudy (anonymous):

ok so am there 2 but there little and i been having trouble with them a) simplify 12+3(2y-3) b) factor the expression completely 18b - 12

OpenStudy (anonymous):

@suckerofmath

OpenStudy (anonymous):

ok it cuase i didnt know if u got them by the text so i just put them here

OpenStudy (anonymous):

For that one, you use the distributive property to multiply the number in the ( ) by the number outside of it. So, you would multiply (3)(2y)+(3)(-3). Remember, a positive times a negative is a negative. Then, you would combine like terms.

OpenStudy (anonymous):

ok so it going to be 6 + 9 then or am i wrong ??

OpenStudy (anonymous):

Close. You have a (y) in the equation and you're multiplying a negative by a positive which gives you a negative number.

OpenStudy (anonymous):

so it gonna be -6 + 9 = 3

OpenStudy (anonymous):

right ?? or na

OpenStudy (anonymous):

Close. You forgot your Y and the 9 should be negative. What's (3) x (2y)?

OpenStudy (anonymous):

it amm -6y ??

OpenStudy (anonymous):

You got the 6y but it should be positive. What's (3) x (-3)?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

??

OpenStudy (anonymous):

It would be 3 x -3. You're answer will be the opposite of 3 x 3 meaning you can multiply 3 x 3 make it negative.

OpenStudy (anonymous):

9 then

OpenStudy (anonymous):

Make it a negative.

OpenStudy (anonymous):

-9

OpenStudy (anonymous):

Yes! Now you should have 12 - 9 + 6y

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Now you just combine like terms

OpenStudy (anonymous):

so now i subtract 12-9 = 3 3+6y =9y

OpenStudy (anonymous):

is it right or not???

OpenStudy (anonymous):

The first thing you should do is 12 - 9

OpenStudy (anonymous):

ok wich is 3

OpenStudy (anonymous):

and then y combine the 3 with the 6y and sums up to 9y no ??

OpenStudy (anonymous):

I'm not sure where you got the 9y, sorry. After the 12 - 9, you should have 3 + 3y

OpenStudy (anonymous):

oo ok ya i got confuse

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

It's alright. What do you think you do next?

OpenStudy (anonymous):

amm ad them up???

OpenStudy (anonymous):

*add

OpenStudy (anonymous):

If both of them had a (y) you would do that. Since they don't, you get them y by it's self. So, you would subtract 3y from both sides, 3 + 3y - 3y = (0 - 3y)

OpenStudy (anonymous):

soo amm it 3y

OpenStudy (anonymous):

I'm not sure where you got that. You should end up with 3 = -3y

OpenStudy (anonymous):

ooo ok sorry

OpenStudy (anonymous):

It's alright. Now you just simplify the equation by dividing both sides by -3 to get -1 = y

OpenStudy (anonymous):

ok so it end up to be -1y